# Retro: Reusing teacher projection head for efficient embedding distillation on Lightweight Models via Self-supervised Learning

### v1

@@ 4.3 Semi-supervised Linear Evaluation
R-152 SEED [9] 54.4 80.1 59.9 38.4 57.0 41.0 33.3 53.7 35.3
DisCo [10] 56.6 80.8 63.4 39.4 58.7 42.7 34.4 55.4 36.7
BINGO [27] - - - 39.9 59.4 43.5 35.7 56.5 38.2
@@ 4.4 Transfer to CIFAR-10/CIFAR-100
We conducted further evaluations to assess the generalization of representations obtained by Retro on CIFAR-10 and CIFAR-100 datasets, using ResNet-18/EfficientNet-B0 as a student and ResNet-50/ResNet-101/ResNet-152 as a teacher. The models were fine-tuned for $100$ epochs, with an initial learning rate of $3$ and the learning rate scheduler decreasing by a factor of $10$ at $60$ and $80$ epochs. All images were resized to $224\times 224$, following the methodology from [9] since the original image resolution of the CIFAR dataset is $32\times 32$. The results presented in Figure 4 show that Retro outperforms prior methods SEED and DisCo across the datasets. Furthermore, the improvement brought by Retro becomes more apparent as the quality of the teacher models improves.
@@
Figure 4: Top-1 accuracy for transfer learning on CIFAR-10 (4a, 4b) and CIFAR-100 (4c, 4d) dataset.
@@ 4.5 Transfer to Detection and Segmentation
For segmentation and downstream detection tasks, we adopt the fine-tuning methodology used in SEED [9] and DisCo [10], where all parameters are fine-tuned. For the detection task on VOC [8], the model is initialized with an initial learning rate of $0.1$, with $200$ warm-up iterations, and the learning rate is decreased by a factor of $10$ at $18$k and $22.2$k steps. The detector is trained for $48k$ steps, with a total batch size of $32$. During training, the input is randomly scaled from $400$ to $800$, and set to $800$ during inference. For detection and segmentation on COCO [17], the model is trained for $180k$ iterations with an initial learning rate of $0.11$. During training, the image scale is randomized from $600$ to $800$.
We also performed tests on detection and segmentation tasks for generalization analysis. Faster R-CNN [22] based on C4 is used for object detection for VOC and R-CNN Mask [13] is used for object detection and version segmentation for COCO. The results are displayed in Table 3. In object detection, our method can yield clear improvements for both VOC and COCO datasets. Also, as claimed by SEED [9], the COCO training dataset has 118k images, while VOC has only 16.5k training images, so the improvement of COCO is relatively small compared to VOC. Therefore, the gain from initialization weight is relatively small. Retro also has an advantage when it comes to instance segmentation tasks.
4.6 Impact of each contribution:
@@
4.7 Computational Complexity:
As illustrated in Figure 3, the computational cost of Retro is higher compared to SEED [9] and DisCo [10] due to the additional forward propagation required for the mean student. The total number of forward propagation is $6$, which is three times higher than SEED and MoCo-V2. However, these additional forward propagations are not used during inference, so there is no overhead at inference time. The results in Table 5 show that Retro has a lower number of learnable parameters than DisCo. Therefore, the run-time overhead of Retro is small and negligible compared to DisCo and BINGO. It should be noted that BINGO requires a KNN run to create a bag of positive samples, while Retro is an end-to-end approach.
@@ 4.7 Computational Complexity:
4.8 Comparison with other Distillation:
@@
Method Top-1
MoCo-V2 (Baseline) [14] 52.2
MoCo-V2 + KD [9] 55.3
MoCo-V2 + RKD [20] 61.6
DisCo + KD [10] 60.6
DisCo + RKD [10] 60.6
BINGO [27] 61.4
Retro 62.9
For further verifying the strengths of Retro, we conducted the comparison against several different distillation strategies. We include feature-based distillation (KD) and relation-based distillation (RKD), following DisCo [10] and BINGO [27]. As shown in Table 6, Retro shows superior performance compared with other distillation methods and surpasses them by a large margin.
4.9 Adapter settings
@@
5 Frozen vs unfrozen projection head
@@
6 Conclusion
7 Funding
This work was supported by the National Cancer Center Grant (NCC-2311350-2).
8 Ethical Statement
This study does not contain any studies with human or animal subjects performed by any of the authors.
9 Conflicts of Interest
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
10 Data Availability Statement
The datasets ImageNet for this study can be downloaded at: https://www.image-net.org/.

### v3

@@ 4.3 Semi-supervised Linear Evaluation
R-152 SEED Fang et al. (2021) 54.4 80.1 59.9 38.4 57.0 41.0 33.3 53.7 35.3
DisCo Gao et al. (2022) 56.6 80.8 63.4 39.4 58.7 42.7 34.4 55.4 36.7
BINGO Xu et al. (2021) - - - 39.9 59.4 43.5 35.7 56.5 38.2
@@ 4.4 Transfer to CIFAR-10/CIFAR-100
We conducted further evaluations to assess the generalization of representations obtained by Retro on CIFAR-10 and CIFAR-100 datasets, using ResNet-18/EfficientNet-B0 as a student and ResNet-50/ResNet-101/ResNet-152 as a teacher. The models were fine-tuned for $100$ epochs, with an initial learning rate of $3$ and the learning rate scheduler decreasing by a factor of $10$ at $60$ and $80$ epochs. All images were resized to $224\times 224$, following the methodology from Fang et al. (2021) since the original image resolution of the CIFAR dataset is $32\times 32$. The results presented in Figure 4 show that Retro outperforms prior methods SEED and DisCo across the datasets. Furthermore, the improvement brought by Retro becomes more apparent as the quality of the teacher models improves.
@@
Figure 4: Top-1 accuracy on CIFAR-10 (4(a), 4(b)) and CIFAR-100 (4(c), 4(d)) dataset.
@@ 4.5 Transfer to Detection and Segmentation
For segmentation and downstream detection tasks, we adopt the fine-tuning methodology used in SEED Fang et al. (2021) and DisCo Gao et al. (2022), where all parameters are fine-tuned. For the detection task on VOC Everingham et al. (2015), the model is initialized with an initial learning rate of $0.1$, with $200$ warm-up iterations, and the learning rate is decreased by a factor of $10$ at $18$k and $22.2$k steps. The detector is trained for $48k$ steps, with a total batch size of $32$. During training, the input is randomly scaled from $400$ to $800$, and set to $800$ during inference. For detection and segmentation on COCO Lin et al. (2014), the model is trained for $180k$ iterations with an initial learning rate of $0.11$. During training, the image scale is randomized from $600$ to $800$.
We also performed tests on detection and segmentation tasks for generalization analysis. Faster R-CNN Ren et al. (2015) based on C4 is used for object detection for VOC and R-CNN Mask He et al. (2017) is used for object detection and version segmentation for COCO. The results are displayed in Table 3. In object detection, our method can yield clear improvements for both VOC and COCO datasets. Also, as claimed by SEED Fang et al. (2021), the COCO training dataset has 118k images, while VOC has only 16.5k training images, so the improvement of COCO is relatively small compared to VOC. Therefore, the gain from initialization weight is relatively small. Retro also has an advantage when it comes to instance segmentation tasks.
5 Conclusion
Appendix A Impact of each contribution:
@@
Appendix B Computational Complexity:
As illustrated in Figure 3, the computational cost of Retro is higher compared to SEED Fang et al. (2021) and DisCo Gao et al. (2022) due to the additional forward propagation required for the mean student. The total number of forward propagation is $6$, which is three times higher than SEED and MoCo-V2. However, these additional forward propagations are not used during inference, so there is no overhead at inference time. The results in Table 5 show that Retro has a lower number of learnable parameters than DisCo. Therefore, the run-time overhead of Retro is small and negligible compared to DisCo and BINGO. It should be noted that BINGO requires a KNN run to create a bag of positive samples, while Retro is an end-to-end approach.
@@ Appendix B Computational Complexity:
Appendix C Comparison with other Distillation:
Method Top-1
MoCo-V2 (Baseline) He et al. (2020) 52.2
MoCo-V2 + KD Fang et al. (2021) 55.3
MoCo-V2 + RKD Park et al. (2019) 61.6
DisCo + KD Gao et al. (2022) 60.6
DisCo + RKD Gao et al. (2022) 60.6
BINGO Xu et al. (2021) 61.4
Retro 62.9
@@ Appendix C Comparison with other Distillation:
For further verifying the strengths of Retro, we conducted the comparison against several different distillation strategies. We include feature-based distillation (KD) and relation-based distillation (RKD), following DisCo Gao et al. (2022) and BINGO Xu et al. (2021). As shown in Table 6, Retro shows superior performance compared with other distillation methods and surpasses them by a large margin.
Appendix D Adapter settings
@@
Appendix E Frozen vs unfrozen projection head
@@
Table 7: Comparison for the training scheme of Retro.